OpenStudy (anonymous):

answer correctly win a medal 1/2x + 1/3y = 3 Which of the following equations is equivalent to the given equation? 3x + 2y = 3 2x + 3y = 18 3x + 2y = 18

OpenStudy (anonymous):

\[6(\frac{ 1 }{ 2 }x + \frac{ 1 }{ 3 }y = 3)\] Multiply everything by 6 to get rid of fractions. 3x + 2y = 18